# Adding numbers with spiking neural circuits on neuromorphic hardware

### v1

@@ Abstract
Progress in neuromorphic computing requires efficient implementation of standard computational problems, like adding numbers. Here we implement one sequential and two parallel binary adders in the Lava software framework, and deploy them to the neuromorphic chip Loihi 2. We describe the time complexity, neuron and synaptic resources, as well as constraints on the bit width of the numbers that can be added with the current implementations. Further, we measure the time required for the addition operation on-chip. Importantly, we encounter trade-offs in terms of time complexity and required chip resources for the three considered adders. While sequential adders have linear time complexity $\mathcal{O}(n)$ and require a linearly increasing number of neurons and synapses with number of bits $n$, the parallel adders have constant time complexity $\mathcal{O}(1)$ and also require a linearly increasing number of neurons, but nonlinearly increasing synaptic resources (scaling with $n^{2}$ or $n\sqrt{n}$). This trade-off between compute time and chip resources may inform decisions in application development, and the implementations we provide may serve as a building block for further progress towards efficient neuromorphic algorithms.
Index Terms:
neuromorphic algorithms, Loihi 2, binary addition, spiking neural networks
#co-senior authors
@@ I Introduction
Moving beyond the standard von Neumann architecture, neuromorphic hardware enables highly parallel, energy-efficient (spike-based) computing [1, 2]. Nevertheless, to make use of these capabilities via next generation algorithms, a variety of reusable software building blocks need to be developed. Through software libraries like Lava [3], the foundation has been laid for modularized implementations of neuromorphic algorithms.
While the promise of neuromorphic hardware lies in neuro-inspired, highly parallel representations and computation, there is also need for interfacing neuromorphic and standard computing systems. Currently, one main bottleneck in neuromorphic systems is in transferring data to and from the system. To deal with these input and output bottlenecks, on-chip implementations of basic functionality, including basic arithmetic operations, are needed.
Addition is one of the fundamental arithmetic operations and is thus required by all modern computing applications. Therefore, to enable the development of new algorithms adapted to neuromorphic hardware, implementations of binary addition in neural circuits of spiking neurons are required. Here, we contribute to the set of computational elements for neuromorphic hardware a set of binary adders, implemented using circuits of spiking neurons in Lava. This enables us to test the adders on the neuromorphic chip Loihi 2 and also provides a basis for future implementations on other neuromorphic systems.
As a major outcome of our investigations, we find that there are trade-offs between the different adders with respect to their time complexity and required number of neurons and synapses. In this context, we characterize the bit precision enabled by the different algorithms and demonstrate under which circumstances an algorithm makes efficient use of compute resources.
@@ II Background
Motivated by the progress in the field of neuromorphic computing, recent research has explored integer adders specifically designed for current neuromorphic hardware systems. Binary encoding of integers for addition has been shown to be efficient and balanced regarding time, neurons, synapses and energy consumption when compared to binning, rate-based or time-based encodings [7]. Recently, multiple binary adder designs have been created and tested on neuromorphic hardware emulators [7, 8] and on hardware [9, 10]. These designs enable efficient use of neurons and synapses under specific circumstances. Furthermore, Date et al. [7] extended simple binary encoding to include additional bits for negative and rational parts, and Wurm et al. [8] used the established two’s complement method to encode signed integers. However, each of these solutions operate in a sequential manner, meaning that the number of required time steps depends on the number of bits that are considered. To the best of our knowledge, up to now, a neuromorphic implementation of parallel adders as proposed by [5] has not been reported.
Aimone et. al implemented a streaming adder that processes one bit at a time using the same neurons and synapses, and recurrent connections for the carry signal [9]. This leads to a resource efficient implementation when the streaming format is applicable. If the sum has to be represented using multiple neurons at the same time step, converting the stream to that format requires further resources.
@@ III Methods
The sum bit is $1$ if (i) any of the three input bits is $1$ and the carry out is 00, or (ii) if all three inputs and the carry out are $1$. This is implemented using a LIF neuron with threshold $1$ and weighted synaptic connections with weight $1$ for the input bits $X_{i}$, $Y_{i}$ and the carry in bit $C_{i-1}$ and weight $-2$ for the carry out bit.
@@ III-A Sequential adder
As the carry is computed with a single neuron and is propagated to the next carry neuron immediately, the adder requires one additional time step per bit, and one further step to compute all the sums. To allow input and output bits to be synchronized, the implementation of the sequential adder makes use of synaptic delays. As Loihi 2 has a maximum synaptic delay of 62 time steps, and the input spikes are delayed to the sum neurons by exactly $n$ steps, this adder only works up to $62$ bits of precision. If higher precision is required, this could be circumvented by employing further neurons to relay the signal, increasing the delay and maximum precision by $64$ bits.
@@ III-B DCTA2 parallel adder
Compared to the sequential adder that only uses weights of $1$ and $-2$, DCTA2s synaptic weights grow exponentially with the number of bits to be added. Our implementation supports DCTA2 with up to 16 bits of precision, limited by these exponential weights. A single synapse group on Loihi allows for 8 bits of precision in the weight mantissas. To achieve 16 bits, we add a second synapse group with the weight exponent increased by 8, allowing for another 8 bits of weight precision. A third synapse group cannot be used for a further increase of precision because this would exceed the limit of the weight exponent allowed by Loihi 2.
@@ IV Results
We implemented each of the described binary adders as a SubProcess in the Lava framework and deployed them to an OheoGulch Loihi 2 system (with chip version N3B3 and Lava-Loihi v0.7.0). We first provide an overview of the time complexity, the required neurons and synapses, as well as the restrictions on the bit precision of numbers that can be processed by each adder (Table I). Then, by adding numbers on Loihi 2 with each of the implemented binary adders, we demonstrate how the runtime scales with the number of bits in each to-be-added number (Figs. 4 and 5).
@@
[7, 8]a $n+1$ $14n-$ 1 $612n-$ 6 $63$
streaming [9] $n+1$ $4$ $9$infty
sequential $n+1$ $2n$ $27n-$ 2 $62$
DCTA2 $2$ $2n$ $n^{2}+5n-1$ $16$
DCTA3 $3$ $4n$ $3n\sqrt{n}+7n-1^{\mathrm{b}}$ $42$
@@
The adders in references [7, 8], which we consider for comparison, have solved slightly different problems. They allow for adding integers of different bit length, have a separate adder for adding positive and negative parts, or include an additional bit in the output to account for potential overflow. For the comparison in Table I we removed all neurons and synapses that were not necessary for adding two numbers of $n$ bits with an $n$ bit result.
@@
V Conclusion
In this work we described three designs for spike-based binary adders, and further implemented and tested them on Loihi 2 using the Lava framework. The implementations are readily available to be integrated as components into larger neuromorphic applications.
All three designs have their use cases. DCTA2 computes addition in the fastest manner, while using many synapses and having limited bit precision due to exponential synaptic weights. The sequential adder supports high-precision addition using fewer synapses, at the cost of runtime growing linearly with the number of bits. Compared to other linear time algorithms [7, 8], the sequential adder requires fewer resources, both in terms of neurons and synapses. DCTA3 provides a balance between DCTA2 and sequential adder regarding runtime, available precision, and synapse usage, at the cost of doubling the number of neurons.
@@ V Conclusion
Other population codes for encoding integers in spiking neural networks offer alternatives to binary encoding. However, they scale linearly with the largest representable number, either in time or neuron count, whereas binary encoding scales logarithmically [7]. On the other hand binary representations can be susceptible to bit flip errors, which may arise in noisy or analog settings, in which case other representations may be more robust.
Using the specific capabilities of the Loihi 2 chip may allow further improvement of the algorithm implementations. In particular, as mentioned above, custom neuron models defined via Loihi 2’s microcode could serve to improve the performance, in terms of neuron and synapse numbers as well as runtime. Graded spikes supported by Loihi 2 could also be used to reduce computation and communication between neurons. However, such adjustments will yield less general formulations of the neuromorphic algorithms compared to those we have provided.
Through their parallel, efficient architecture, neuromorphic hardware systems offer a promising platform for algorithms in general, even beyond cognitive/biologically-constrained implementations [13]. Given the bottleneck in transferring data between neuromorphic chips and standard computers, solutions are required to minimize data movement. A mixture of neuro-inspired algorithms and on-chip implementations of standard tasks like data processing, analysis, and optimization, while perhaps computationally less efficient in time per operation, can minimize data movement and improve overall efficiency. Adder circuits constitute a critical piece in this toolbox, which we have addressed by characterizing three optimized, generally applicable neuromorphic adders.

### v2

@@ Abstract
[…]
#co-senior authors  *corresponding author
@@ I Introduction
The promise of neuromorphic hardware lies in highly parallel and energy-efficient representations and computation, taking inspiration from networks of neurons in the brain [1, 2]. However, not all tasks lend themselves to decentralized, parallelized implementations, meaning that standard computation will be required in conjunction with – and not replaced by – neuro-inspired neuromorphic algorithms. Future computing systems will thus often be of hybrid nature, depending on the integration of CPUs and GPUs, related memory systems, as well as novel decentralized chip architectures like neuromorphic hardware.
As for any computing architecture, transferring data between systems is a bottleneck, which also holds true for neuromorphic chips. This raises a critical question: Should a system transfer data from a neuromorphic chip to standard hardware to perform certain computations and then transfer this processed data back to the neuromorphic chip? Or would it overall be more efficient to develop decentralized implementations of standard functionality to be used on the neuromorphic system, in order to avoid the data transfer bottleneck (despite the implementation itself potentially being slower and/or less efficient)?
The answer to this question ultimately depends on the details of the considered system. However, when the amount of data to transfer becomes large enough, an on-chip implementation of standard computational operations is likely to outperform solutions requiring data transfer. For example, in a large-scale neuro-inspired architecture, it may be more prudent to implement a cost function for optimization, or a simple procedure for data analysis, using arithmetic operations directly on chip instead of transferring data back and forth to and from a CPU.
Given this issue, we set out towards on-chip implementations of standard operations with the goal of minimizing unnecessary data transfer. To start in this direction, here we consider addition, one of the fundamental arithmetic operations required by all modern computing applications. Addition is needed for example when computing error functions for online learning [3, 4, 5, 6], or when calculating shortest paths in graph search algorithms that are optimally suited for parallelized implementation on neuromorphic hardware [7].
Here, we contribute to the set of computational elements for neuromorphic hardware a set of binary adders, implemented using circuits of spiking neurons in Lava. This enables us to test the adders on the neuromorphic chip Loihi 2 and to provide a basis for future implementations on other neuromorphic systems. As a major outcome of our investigations, we find that there are trade-offs between the different adders with respect to their time complexity and required number of neurons and synapses. In this context, we characterize the bit precision enabled by the different algorithms and demonstrate under which circumstances an algorithm makes efficient use of compute resources.
@@ II Background
Motivated by the progress in the field of neuromorphic computing, recent research has explored integer adders specifically designed for current neuromorphic hardware systems. Binary encoding of integers for addition has been shown to be efficient and balanced regarding time, neurons, synapses and energy consumption when compared to binning, rate-based or time-based encodings [11]. Recently, multiple binary adder designs have been created and tested on neuromorphic hardware emulators [11, 12] and on hardware [13, 14, 15]. These designs enable efficient use of neurons and synapses under specific circumstances. Furthermore, Date et al. [11] extended simple binary encoding to include additional bits for negative and rational parts, and Wurm et al. [12] used the established two’s complement method to encode signed integers. However, each of these solutions operate in a sequential manner, meaning that the number of required time steps depends on the number of bits that are considered. To the best of our knowledge, up to now, a neuromorphic implementation of parallel adders as proposed by [9] has not been reported.
Aimone et al. implemented a streaming adder that processes one bit at a time using the same neurons and synapses, and recurrent connections for the carry signal [13]. This leads to a resource efficient implementation when the streaming format is applicable. If the sum has to be represented using multiple neurons at the same time step, converting the stream to that format requires further resources.
@@ III Methods
The sum bit is $1$ if (i) any of the three input bits is $1$ and the carry out is 0, or (ii) if all three inputs and the carry out are $1$. This is implemented using a LIF neuron with threshold $1$ and weighted synaptic connections with weight $1$ for the input bits $X_{i}$, $Y_{i}$ and the carry in bit $C_{i-1}$ and weight 2-2 for the carry out bit.
@@ III-A Sequential adder
As the carry is computed with a single neuron and is propagated to the next carry neuron immediately, the adder requires one additional time step per bit, and one further step to compute all the sums. To allow input and output bits to be synchronized, the implementation of the sequential adder makes use of synaptic delays. As Loihi 2 has a maximum synaptic delay of 63 time steps [19], and the input spikes reach the summing neurons with a delay of exactly $n$ steps, this adder only works up to $63$ bits of precision. If higher precision is required, a relay neuron with threshold 0 can be inserted in place of each synapse. This adds one time step of delay for the processing in the neuron and up to another $63$ time steps of delay through the second synapse, therefore increasing the delay and maximum precision by $64$ bits for each added layer of relay neurons.
@@ III-B DCTA2 parallel adder
Compared to the sequential adder that only uses weights of $1$ and 2-2, DCTA2s synaptic weights grow exponentially with the number of bits to be added. Our implementation supports DCTA2 with up to 16 bits of precision, limited by these exponential weights. A single synapse group on Loihi allows for 8 bits of precision in the weight mantissas. To achieve 16 bits, we add a second synapse group with the weight exponent increased by 8, allowing for another 8 bits of weight precision. A third synapse group cannot be used for a further increase of precision because this would exceed the limit of the weight exponent allowed by Loihi 2.
@@ IV Results
We implemented each of the described binary adders as a SubProcess in the Lava framework and deployed them to an OheoGulch Loihi 2 system (with chip version N3B3 and Lava-Loihi v0.7.0). We first provide an overview of the time complexity, the required neurons and synapses, as well as the restrictions on the bit precision of numbers that can be processed by each adder (Table I). Then, by adding numbers on Loihi 2 with each of the implemented binary adders, we demonstrate how the runtime scales with the number of bits in each to-be-added number (Figs. 4 and 7).
@@
VN[11, 12]a $n+1$ 4n14n-1 12n612n-6 $63$
streaming [13] $n+1$ 44 99 $\infty$
sequential $n+1$ $2n$ 7n27n-2 $62$
DCTA2 22 $2n$ $n^{2}+5n-1$ $16$
DCTA3 33 $4n$ $3n\sqrt{n}+7n-1^{\mathrm{b}}$ $42$
@@
The ‘virtual neuron’ (VN) adder from references [11, 12], which we consider for comparison, has solved a slightly different problems. They allow for adding integers of different bit length, have a separate adder for adding positive and negative parts, or include an additional bit in the output to account for potential overflow. For the comparison in Table I we removed all neurons and synapses that were not necessary for adding two numbers of $n$ bits with an $n$ bit result. The streaming adder [13] uses a different format for input and output than the other four adders. Inputs are read, starting with the least significant bit, one bit in each timestep through the same synapse and output is produced in the same way. This allows for reusing neurons for the computation across bits, leading to a circuit size that is independent of the number of bits in the input. While the table gives theoretical neuron and synapse counts, actual resource usage may be different, depending on the hardware platform. For example, on Loihi 2, the number of available neurons is reduced if longer delays are used [19]. We measured the actual resource usage for our three target adders, as well as for an implementation of the VN adder, on Loihi 2 and found that every one of our target adders, within its supported range, fits on a single neuromorphic core (Fig. 5). This comparison reveals that DCTA3 uses less resources on Loihi 2 than the sequential adder because of lower delays, even though the theoretical metrics in Table I suggest that DCTA3 uses more neurons and synapses. When comparing the sequential adder to the other linear time adder VN, we can see that the theoretical reduction in synapses and neurons translates directly to a reduction in resource usage on the Loihi 2 hardware (Fig. 5), as both designs utilize similar synaptic delays to synchronize inputs and outputs.
@@
Figure 5: Same experiments as in Fig. 4. Chip resource usage as measured by Lava, as a fraction of a single neuromorphic core. This includes neurons for providing input and neurons for reading output. The jumps in neuron usage occur at powers of 2, where an additional bit for the maximum synaptic delays halves the neuron capacity on Loihi 2.
Figure 6: Spikes and synaptic events measured using a Lava Profiler. Input and output spikes are included. For the sequential adder and DCTA2, the two measures are identical, but for DCTA3 there are more synaptic events than spikes.
@@
V Discussion
In this work we have described three designs for spike-based binary adders, and further implemented and tested them on Loihi 2 using the Lava framework. The implementations are readily available to be integrated as components into larger neuromorphic applications.
All three designs have their use cases. DCTA2 computes addition in the fastest manner, while using many synapses and having limited bit precision due to exponential synaptic weights. The sequential adder supports high-precision addition using fewer spikes and synaptic operations, being more efficient in theory. This comes at the cost of runtime growing linearly with the number of bits. Further, in the specific case of Loihi 2, the theoretical advantage in terms of synapse count is outweighed by the reduced resources available due to the necessary use of high delays. Compared to the other linear time algorithm [11, 12], which use the same delays in the implementation, our sequential adder requires fewer resources, both in terms of neurons and synapses. DCTA3 provides a balance between DCTA2 and sequential adder regarding runtime, available precision, and synapse usage, at the cost of doubling the number of neurons. Depending on the specific application and neuromorphic platform, the trade-off between runtime, hardware resources and energy consumption might be different. Energy consumption may be estimated based on chip characteristics and the numbers of spikes and synaptic operations reported in this work.
@@ V Discussion
Other population codes for encoding integers in spiking neural networks offer alternatives to binary encoding. Binning assigns a neuron to each possible value in the representation, increasing the required space exponentially. Rate-based representations require many spikes to represent a single number, resulting in exponential time. With the same time requirements, single-spike-timing-based representation uses fewer spikes, thus reducing the energy required compared to the rate-based approach. While these representations are less efficient in time and use of neurons compared to binary representation, they offer other benefits such as increased robustness to errors in the input. In particular, a fault in the high bits can significantly change the interpretation of a binary number, which may be prevented by a rate-based representation. Moreover, when only a few operations are required and further processing before and after the addition requires another representation, it may be beneficial to use that representation for the addition and avoid the conversion.
The specific implementation provided is aimed at neuromorphic hardware systems with a fixed, not fully programmable, neuron model such as Loihi 1 (digital) or BrainScales-2 (analog), which can under certain circumstances enable extremely efficient emulation of spiking networks [20, 21]. These systems require neuron-based algorithms to increase the scope of problems they can solve. On fully programmable hardware like Loihi 2 or SpiNNaker 2, making direct use of the underlying processing architecture without adding the abstraction of neurons and synapses will be more efficient for standard arithmetic. However, as stated above, the minimal approach of hardware with fixed neuron models is promising, and therefore we aim to enhance its usability by implementing neuron-based arithmetic primitives. This vision is shared by others, constructing arithmetic logic units (ALUs) with multiple operations in a similar manner [15].
Through their parallel, efficient architecture, neuromorphic hardware systems offer a promising platform for algorithms in general, even beyond cognitive/biologically-constrained implementations [22, 23]. Given the bottleneck in transferring data between neuromorphic chips and standard computers, solutions are required for on-chip computation to minimize data movement. While data transfer rates and latency between neuromorphic systems will improve in the future, data transfer will remain a bottleneck compared to the increasing speed of computation. A mixture of neuro-inspired algorithms and on-chip implementations of standard tasks like data processing, analysis, and optimization, although perhaps computationally less efficient in time per operation, can minimize data movement and improve overall efficiency.
To achieve such implementations in spiking networks, multiple building blocks are required – from basic logic gates [15] to arithmetic operations such as addition and multiplication [13, 24]. Implementations of such building blocks can take inspiration from neuroscience [25], threshold logic [10], and traditional digital circuits. Adder circuits constitute a critical piece in this toolbox, which we have addressed by characterizing a set of optimized, generally applicable neuromorphic adders.